The larger of tow consecutive integers is 10 greater than twice the smaller. Find the integers.
Answer by edjones(7311)

Let the integers be x and x+1
2x+10=x+1
subtract x and 10 from both sides
x=-9 x+1=-8
Check:
-8=(2*-9)+10
-8=-18+10
Ed